What percentage of 2, 4-dinitrophenol (pK_a =4.1)
I need the answer to this question as well work shown.
Answer #1

14.

pH = 8.2

pKa = 4.1

Using,

pH = pKa + log(deprotonated/protonated)

8.2 = 4.1 + log(deprotonated/protonated)

(deprotonated/protonated) = 12589.254

percentage of ionization = 12589.254 x 100/12590.254 = 99.99%